In the endothermic process the volume of a gas contracted from 180 ml to 175ml against a constant pressure of 722 during the process 30 cal of heat energy were given off what was the internal energy change for the system in joules


1
Expert's answer
2021-12-14T09:27:32-0500

"30\ cal=125.52\ J"


"Q=\Delta U+p\Delta V \to \Delta U=Q-p\Delta V="


"=-125.52+722000\cdot0.000005=-121.91\ (J)" . Answer